OpenStudy (anonymous):

Evaluate the following expression using the values given: Find 5x - 3y - z if x = -2, y = 2, and z = -3 -19 19 13 -13

OpenStudy (anonymous):

Plug it in.

OpenStudy (anonymous):

could ya help me with it?

OpenStudy (anonymous):

@curiouscrab

OpenStudy (anonymous):

Replace x with the x value, y with the y value, and z with the z value. So, 5(-2)-3(2)-(-3)

OpenStudy (anonymous):

Then, simplify. -10-6+3

OpenStudy (anonymous):

-10-6=-16 -16+3=-13

OpenStudy (anonymous):

so -13?

OpenStudy (anonymous):

Yes.